# Bilayer graphene in periodic and quasiperiodic magnetic superlattices

### v1

@@ 5 Conclusions
Starting from the tight binding model for the low-lying excitations of bilayer graphene, we have determined the energy band structure of bilayer graphene when a periodic magnetic field is applied. In this approach, which generalizes previous works about monolayer graphene, we have employed second order SUSY transformations to deal with the topological features of electronic states in bilayer graphene. The bound states created by the transformations are often related to impurities or van Hove singularities. Alike to the case of monolayer graphene, the implementation of the most general Darboux transformation breaks the periodicity of the originally periodic magnetic field, due to the introduction of such bound states. However, afar from these introduced defects in graphene the solutions turn out to be periodic, behaving as it was expected. It is worth to note that the number of experimental studies about electronic transitions have increased in the last years, with its optimization being a major concern. About this point, the present study mimics the effect of gold island-enhanced multiplex quantum dots, in which the quantum dots act as a bridge between the gold islands. On the other hand, the present analysis reinforces the role of the second order supersymmetric quantum mechanics for studying electronic properties of bilayer graphene. Throughout this work we restricted ourselves to the simplest case of the Lamé potentials; nevertheless, we might choose, between a long list of candidates, another examples to illustrate the effects of the Darboux transformations for bilayer graphene.
@@ Acknowledgments
The authors acknowledge the support of Consejo Nacional de Humanidades Ciencia y Tecnología (CONAHCyT-México) under grant FORDECYT-PRONACES/61533/2020. OPT acknowledges
CONAHCyT by a postdoctoral fellowship.

### v2

@@ 5 Conclusions
Van der Waals heterostructures, such as bilayer graphene, offer an alternative amidst the advent and rapid progress of nanotechnology. Their development has increased the availability of atomic-layer materials, encompassing semiconductors, semimetals, superconductors, ferromagnets, and topological insulators. This progress has parallel advancements in theoretical frameworks, facilitating the rapid creation of quantum metamaterials for electrons. The stacking of two-dimensional layers atop one another forms a superlattice, where electrons can be described by Bloch bands within an enlarged unit cell. These lattices feature a corresponding tunable Brillouin zone, enabling precise control over the arrangement of electron and hole filling in the heterostructure. The ability to engineer unit cell sizes and electron filling can yield significant outcomes, particularly when applying a magnetic field. Quantum metamaterials aim to manipulate and control bound states embedded in superlattices and electromagnetic waves at the quantum level, enabling functionalities such as tunable optical properties, efficient light-matter interactions and novel quantum devices like quantum sensors or transistors. In this direction, starting from the tight binding model for the low-lying excitations of bilayer graphene, we have determined the energy band structure of bilayer graphene when a periodic magnetic field is applied. In this approach, which generalizes previous works about monolayer graphene, we have employed second order SUSY transformations to deal with the topological features of electronic states in bilayer graphene. The bound states created by the transformations are often related to impurities or van Hove singularities. Alike to the case of monolayer graphene, the implementation of the most general Darboux transformation breaks the periodicity of the originally periodic magnetic field, due to the introduction of such bound states. However, afar from these introduced defects in graphene the solutions turn out to be periodic, behaving as it was expected. It is worth to note that the number of experimental studies about electronic transitions have increased in the last years, with its optimization being a major concern. About this point, the present study mimics the effect of gold island-enhanced multiplex quantum dots, in which the quantum dots act as a bridge between the gold islands. On the other hand, the present analysis reinforces the role of the second order supersymmetric quantum mechanics for studying electronic properties of bilayer graphene. Throughout this work we restricted ourselves to the simplest case of the Lamé potentials; nevertheless, we might choose, between a long list of candidates, another examples to illustrate the effects of the Darboux transformations for bilayer graphene.
@@ Acknowledgments
The authors acknowledge the support of Consejo Nacional de Humanidades Ciencia y Tecnología (CONAHCyT-México) under grant FORDECYT-PRONACES/61533/2020. OPT acknowledges CONAHCyT by a postdoctoral fellowship.